OpenStudy (anonymous):

If sin theta = 2/3 (and it does believe me) and theta is in II (I think that means quadrant two?) find sin 2 theta.

OpenStudy (anonymous):

first you need to find \(\cos(\theta)\) by pythgoras

OpenStudy (anonymous):

then use \[(\sin(2\theta)=2\sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

Wait so you're saying I should draw a triangle first.

OpenStudy (anonymous):

\[\sin(\theta)=\frac{2}{3}\implies\cos(\theta)=-\frac{\sqrt{5}}{3}\] the minus sign because you are in quadrant II

OpenStudy (anonymous):



OpenStudy (anonymous):

then plug in to formula above and you are done

OpenStudy (anonymous):

you said it was that is how

OpenStudy (anonymous):

Haha I know. I just realized.

OpenStudy (anonymous):

If sin theta = 2/3 (and it does believe me)

OpenStudy (anonymous):

so i went from there notice that we are not required to find \(\theta\)

OpenStudy (anonymous):

Now I have a question. This is quadrant two right?

OpenStudy (anonymous):

yes. and there sine is positive (because you are up) and cosine is negative (because you are to the left)

OpenStudy (anonymous):

Now hold on a second what does cos have to do with anything?

OpenStudy (anonymous):

You're saying I should just multiply that cosine stuff with two?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[-2\times \frac{2}{3}\times \frac{\sqrt{5}}{3}\]

OpenStudy (anonymous):

or if you prefer \[-\frac{4\sqrt{5}}{9}\]